OpenStudy (eaglecore):

Establish the identity. (cot u + csc u -1) __________________ = csc u + cot u (cot u - csc u +1) I've gotten it to: ((cos u - sin u)+1) __________________ ((cos u + sin u) -1) That is where I get stuck. :/

OpenStudy (anonymous):

Do you want me to proceed from the actual question or the step where you have stuck now, by the way both are one and the same thing..

OpenStudy (anonymous):

The technique is to use Rationalization Method here.. It means Multiplying and Dividing by the Denominator by changing a sign to simplify the expression easily, as for example here: Denominator is : \(cot(u) - cosec(u) + 1\), so you should multiply and divide by \(cot(u) - cosec(u) \color{red}{-} 1\) (See, I have changed +1 to -1.. Now going further: \[\frac{\cot(u) + cosec(u) - 1}{\cot(u) - cosec(u) +1} \times \frac{\cot(u) - cosec(u) - 1}{\cot(u) - cosec(u) -1}\] \[\implies \frac{[(\cot(u) - 1)^2 - cosec^2(u)]}{[(\cot(u) - cosec(u))^2 - 1]}\] \[\implies \frac{\cot^2(u) + 1 - 2\cot(u) + cosec^2(u)}{\cot^2(u) + cosec^2(u) - 2\cot(u)cosec(u) - 1}\] \[\implies \frac{\cancel{cosec^2(u)} - 2 \cot(u) - \cancel{cosec^2(u)}}{\cot^2(u) + \cancel{1} + \cot^2(u) - 2\cot(u)cosec(u) - \cancel{1}}\] \[\implies \frac{-2 \cot(u)}{2 \cot^2(u) - 2 \cot(u) cosec(u)} \implies \frac{\cancel{-2 \cot(u)}}{\cancel{-2\cot(u)}[cosec(u) - \cot(u)]}\] \[\color{green}{\implies \frac{1}{cosec(u) - \cot(u)}}\]

OpenStudy (anonymous):

now, you are almost there in front of answer, just Rationalize it once more with \(cosec(u) + cot(u)\), just one step you are away from getting there.. :)

OpenStudy (anonymous):

Identities that I have used here is/are: \[a) \color{red}{\; \; 1 + \cot^2(u) = cosec^2(u)}\] \[b) \; \; \color{green}{cosec^2(u) - \cot^2(u) = 1}\] And these are one and the same thing..

OpenStudy (anonymous):

A similar type of question you will find at : http://openstudy.com/study#/updates/516cef8fe4b063fcbbbcbf77 You must see it and try to solve it yourself...

OpenStudy (eaglecore):

Thank you very much! I have already solved this problem at this point, I forgot to come back here and take this down. But your explanation of what is going on helps SO much! Thank you again for taking the time to explain this. :)

OpenStudy (anonymous):

It is okay @EagleCore... You are welcome dear..